OpenStudy (anonymous):

4.06]Which equation represents the line passing through the points (1, 4) and (−2, 13)? 3x + y = 7 3x − y = 7 x − 3y = 7 x + 3y = 7

OpenStudy (anonymous):

Take a general equation of a line: y=ax+b and substiute x and y with respectively x and y coordinates of your two points - you'll have a system of two equations and when you solve them, you will get your a and b.

OpenStudy (anonymous):

they don't give you a graph and i forgot how to do that correctly. it's been a long time since i've done point slope stuff. which i loved

OpenStudy (anonymous):

Ok, so again: you have your line. It passes through those two points. Let's subsitute general x and y by specific x and y coodinates of those two points. y = ax+b gives us: 4 = 1x + b 13 = -2x + b. Do you see where that came from? Then you have two equations and you can solve for a and b, so you'll get numerical values of them.

OpenStudy (anonymous):

If you don't understand, tell me in which part I wasn't clear.

OpenStudy (anonymous):

i get it. so whats the answer? i don't have any graph paper to do this right now

OpenStudy (anonymous):

You don't need to graph anything. Just solve this system of linear equations. Do you know how to do that?

OpenStudy (anonymous):

oh yeah, i hate them though

OpenStudy (anonymous):

ok so i got answer c is that right?

OpenStudy (anonymous):

Well, if you solve them, you get a=-3 (it was supposed to be a, not an x, sorry) and b=7. So y=-3x+7, then y + 3x = 7 and that's the answer.

OpenStudy (anonymous):

ohh i see thanks

OpenStudy (anonymous):

Sure. :)